In the Carco Auto Factory, robots assemble cars. If 3 robots assemble 17 cars in 10 minutes, how many cars can 14 robots assemble in 45 minutes if all robots work at the same rate all the time?
Answer by stanbon(75887)

In the Carco Auto Factory, robots assemble cars. If 3 robots assemble 17 cars in 10 minutes, how many cars can 14 robots assemble in 45 minutes if all robots work at the same rate all the time?
----------------------
# of cars is directly related to # of robots and directly to amt. of time.
----------------------
c = k*r*t
--------
Solve for "k" using "3 robots assemble 17 cars in 10 minutes"
17 = k*3*10
k = 17/30
----
Equation:
c = (17/30)r*t
-----
How many cars can 14 robots assemble in 45 minutes if all robots work at the same rate all the time?
--------
c = (17/30)*14*45
# of cars = 357
==================
Cheers,
Stan H.